OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 2x^2 - 10x + 7 = 0 the answer ive been getting doesn't fit in the answer blank I was given

OpenStudy (anonymous):

Here a = 2, b = -10, c = 7, so ( - b ± √ (b² - 4ac) ) / (2a) = ( 10 ± √ (10² - 4(2)(7)) ) / (2(2)) = ( 10 ± √ (100 - 56) ) / (4) = ( 10 ± √ 44 ) / (4) = ( 5 ± √ 11 ) / (2).

OpenStudy (anonymous):

thank you I was getting 10 (plus or minus) 44

OpenStudy (anonymous):

haha your welcome !